# Bayesian Inference for PDE-based Inverse Problems using the Optimization of a Discrete Loss

### v1

@@ 1 Introduction
Inverse problems are ubiquitous in science, engineering, and medicine, in particular for problems where observations provide only indirect or incomplete information about a system [1]. Inverse problems are central in a wide range of applications such as flow field reconstruction [2, 3, 4], data assimilation [5], medical imaging [6, 7], and parameters estimation of material properties [8, 9, 10]. A particularly challenging class of inverse problems arises when the forward model is governed by ordinary differential equations or partial differential equations [11]. Incorporating physical knowledge through this approach reduces the space of possible solutions, avoiding the need for arbitrary regularization as is often the case in inverse problems [12, 13, 14]. However, this approach can suffer from the high dimensionality of the problem, stiffness, noisy measurements, and sensitivity to parameters. In particular, quantifying the uncertainties of solutions is challenging with standard techniques for inverse PDE problems such as Bayesian inference [15, 14], variational methods [16], ensemble Kalman methods [17], and adjoint-based optimization [18], which can be limited with issues of scalability, robustness, and computational cost.
@@
Despite these advances, PINNs and ODIL solutions can be affected by measurement errors of the provided data. In particular, it is not clear how these measurement errors affect the uncertainties of the solution of the inverse problem. Recent studies have addressed these issues in the context of PINNs [36, 37] through a Bayesian approach to account for the variability of the unknown field. However, a similar theory has not been developed for ODIL. We note that the Bayesian extensions of PINNs have been applied to problems of less than two dimensions. On the other hand, a Bayesian extension of ODIL could potentially provide quantified uncertainties for inverse problems in higher dimensions, as ODIL was shown to be orders of magnitude faster and more robust than PINNs in two and three dimensional benchmark problems [28].
In this study, we present B-ODIL, a Bayesian extension of ODIL. In this framework, the prior incorporates knowledge from the PDE, while the likelihood couples the observed data to the unknown field. This method provides a solution to the inverse problem with quantified uncertainties. We account for computational burden by estimating the posterior distribution with different sampling techniques and approximations. We have designed a series of benchmarks with increasing levels of complexity to test B-ODIL. First, we consider the ODE describing the dynamics of a harmonic oscillator, to compare the validity of the Laplace and mode approximations with the ground truth given by Hamiltonian Monte-Carlo (HMC) sampling in a computationally tractable setting. We then apply B-ODIL to the one dimensional PDE of the diffusion equation with unknown initial conditions, introducing the challenge of ill-posed inverse problems, typical in PDE-based inference, where uncertainty quantification is crucial. The third benchmark tests the method’s ability to reconstruct the states of a non-linear two dimensional PDE from synthetic data, and we demonstrate that the ground truth falls within the uncertainty bounds provided by B-ODIL. Finally, we apply the method to a three dimensional model of tumor growth coupled with real patients data, and provide estimates of tumor cell fields with quantified uncertainties given medical images.
@@ 2.1 ODIL
where $\lambda$ is a positive constant that describes the importance of fitting the data with respect to satisfying the discrete PDE. The minimization of eq. 4 is typically performed with gradient based methods, coupled with auto-differentiation software, simplifying the implementation of this method.
@@ 2.2 B-ODIL: Bayesian inference for inverse problems with ODIL
We now consider the Bayesian formulation of PDE-based inverse problems, where the goal is to infer parameters $\theta$ and the solution field $\mathbf{u}$ from noisy measurements $\mathcal{D}$, given a model in the form of a PDE. In this framework, the posterior distribution is obtained by combining a likelihood model for the data with prior knowledge enforcing compatibility between the solution field and the PDE. As we have seen in the previous section, ODIL provides a natural way to encode this compatibility through a PDE loss. We thus incorporate this PDE loss into the prior information, allowing us to extend ODIL into the Bayesian setting.
We assume that the measurements $\mathcal{D}=\{y_{k}\}_{k=1}^{N_{D}}$ are noisy and that we have a model describing this noise. We would like to estimate the parameters $\mathbf{\theta}$ and the solution $\mathbf{u}$, with quantified uncertainties. According to the Bayes’ theorem, the posterior distribution of these quantities given the data reads
@@
where $\sigma$ is the standard deviation associated with the observable and $h_{k}$ , $k=1,\dots,N_{D}$, are measurement operators on the solution $\mathbf{u}$. We note that several other likelihood functions are compatible with the method presented in this study, and we will give further examples in the results section.
As prior knowledge, we assume that the possible solutions $\mathbf{u}$ are compatible with the PDE associated with the model. To model this, the prior over the solution $\mathbf{u}$ with parameters $\mathbf{\theta}$ is such that the loss $L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})$ is small, and thus we choose priors of the form
$P(\mathbf{u},\mathbf{\theta})=\frac{1}{Z}\exp\left(-\beta L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})\right)P(\mathbf{u})P(\mathbf{\theta}),$ (7)
where $\beta$ is a positive scalar that controls how peaked the distribution is, and $Z$ is a normalization constant that does not depend on $\mathbf{u}$ or $\mathbf{\theta}$. Finally, $P(\mathbf{u})$ and $P(\mathbf{\theta})$ are additional prior knowledge on $\mathbf{u}$ and $\mathbf{\theta}$, respectively. The posterior distribution of the solution and parameters of the model thus becomes
$P(\mathbf{u},\mathbf{\theta}|\mathcal{D})\propto P(\mathcal{D}|\mathbf{u},\mathbf{\theta})\frac{1}{Z}\exp\left(-\beta L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})\right)P(\mathbf{u})P(\mathbf{\theta}).$ (8)
The dimensionality of this problem is large due to the unknown variables $\mathbf{u}$ , challenging sampling techniques. Alternatively, we apply the Laplace approximation to the posterior. The log-posterior reads
$\log P(\mathbf{u},\mathbf{\theta}|\mathcal{D})=\log P(\mathcal{D}|\mathbf{u},\mathbf{\theta})-\beta L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})-\log{Z}-\log P(\mathbf{u})-\log P(\mathbf{\theta})+C,$ (9)
where $C$ is a scalar that does not depend on $\mathbf{\theta}$ or $\mathbf{u}$. We expand this quantity in Taylor series up to second order around the maximum a posteriori (MAP),
@@
where $\mathbf{v}=(\mathbf{u},\mathbf{\theta})$, $\mathbf{v}^{\star}$ is the solution at the MAP, and $H$ is the hessian of the log-posterior evaluated at $\mathbf{v}^{\star}$. Thus, the posterior is approximated with a multivariate Gaussian with mean $\mathbf{v}^{\star}$ and covariance $\Sigma=-H^{-1}$. We note that when the likelihood takes the form of eq. 6, maximizing eq. 9, i.e. finding the MAP, corresponds to the original ODIL method, eq. 4 with $\lambda=N_{D}/\beta$, assuming that the priors $P(\mathbf{u})$ and $P(\mathbf{\theta})$ are uniform with large bounds. Thus, computing uncertainties over predictions that were computed with ODIL consists only in computing the Hessian and inverting it.
@@ 2.3 Inference of model parameters for inverse problems
The evaluation of this high-dimensional integral is intractable. Instead, we assume that the joint distribution is peaked around the MAP $\mathbf{u}^{\star}(\mathbf{\theta})=\operatorname*{arg\,max}\limits_{\mathbf{u}}P(\mathbf{u}|\mathbf{\theta},\mathcal{D})$ and we approximate the posterior distribution of the model parameters using the mode approximation
@@ 3.1 Harmonic oscillator
We first consider a simple example to demonstrate the validity of the Laplace approximation and that of eq. 12. Consider the system of ODEs of a harmonic oscillator with mass $m$ and spring coefficient $k$,
@@
for time $t\in(0,T)$ and with unknown initial conditions. Given measurements of position at known times $\mathcal{D}=\left\{t_{j},x_{j}\right\}_{j=1}^{N_{D}}$ with known precision $\sigma=0.1$, we want to predict the position and velocity of the system throughout the whole time interval. We discretize the position and velocity into $N=64$ equidistant time intervals with $\Delta t=T/N$ and $T=10$. The discrete loss are computed using the midpoint rule,
@@
Finally, the observation functions correspond to linear interpolations of the positions at data points $t_{j}$, $j=1,2,\dots,N_{D}$. For now, we assume that the parameter $\omega=\sqrt{k/m}=1$ is known. We use synthetic data consisting of $N_{D}=20$ time points uniformly distributed in $t\in(0,T/2)$, and corresponding positions with Gaussian distribution centered around the exact solution of the model with standard deviation of $\sigma$. The initial conditions used to generate the data were set to $x(0)=0.5$ and $v(0)=0.2$. We choose a value of $\beta=${10}^{4}\text{\,}$\$ in the prior distribution (see eq. 7). The Laplace approximation of this problem is shown on fig. 1. For comparison, we solve the same problem with HMC, and show the results on the same figure. The HMC results were obtained from 10’000 samples, chain lengths $l=10$, mass matrix $M=1$, and step size $\delta t=0.008$, tuned to for an acceptance rate of $\alpha\approx 0.65$.
Figure 1: Left: Prediction of the position (top row) and velocity (bottom row) of the oscillator given the data (crosses) using the UQ-ODIL framework with Laplace approximation (left column) and HMC (right column). The shaded area denotes the 5 to 95% quantiles of the posterior, and the solid line denotes the posterior mean. The dashed line represents the underlying process that was used to generate the data. Right: Covariance matrices of the full solution $(x_{1},\dots,x_{N},v_{1},\dots,v_{N})$ obtained from the Laplace approximation (left column) and estimated from HMC samples (right column).
@@
$P(\mathbf{\theta}|\mathcal{D})=\frac{(2\pi)^{d/2}}{\det H(\mathbf{\theta})}\exp l(\mathbf{u}^{\star}(\mathbf{\theta}),\mathbf{\theta}),$
@@
where $d$ is the dimension of $\mathbf{u}$. To evaluate this expression for each value of $\theta$, we use the optimal value $\mathbf{u}^{\star}(\mathbf{\theta})$ estimated with ODIL, and use automatic differentiation to compute $H(\mathbf{\theta})$.

### v2

@@ Abstract
keywords:
Inverse problems , Bayesian inference , ODIL , Partial Differential Equations
@@ 1 Introduction
Inverse problems are ubiquitous in science, engineering, and medicine, in particular for problems where observations provide only indirect or incomplete information about a system [20]. Inverse problems are central in a wide range of applications such as flow field reconstruction [21, 15, 29], data assimilation [12], medical imaging [50, 26], and parameters estimation of material properties [4, 17, 2]. A particularly challenging class of inverse problems arises when the forward model is governed by ordinary differential equations or partial differential equations [25]. Incorporating physical knowledge through mathematical models constrains the space of admissible solutions and can reduce the severity of ill-posedness, although inverse problems often remain ill-posed and still require regularization [18, 34, 49]. However, this approach can suffer from the high dimensionality of the problem, stiffness, noisy measurements, and sensitivity to parameters. In particular, quantifying the uncertainties of solutions is challenging with standard techniques for inverse PDE problems such as Bayesian inference [52, 49], variational methods [47], ensemble Kalman methods [24], and adjoint-based optimization [54], which can be limited with issues of scalability, robustness, and computational cost. Bayesian inference for PDE-constrained inverse problems, Laplace and Gauss-Newton-type approximations, has been extensively studied in the context of function-space formulations [49, 10, 45, 5].
@@
Despite these advances, PINNs and ODIL solutions can be affected by measurement errors of the provided data. In particular, it is not clear how these measurement errors affect the uncertainties of the solution of the inverse problem. Recent studies have addressed these issues in the context of PINNs [57, 42] through a Bayesian and a conformal prediction framework [58] approach to account for the variability of the unknown field. However, a similar theory has not been developed for ODIL. We note that the Bayesian extensions of PINNs have been applied to problems of less than two dimensions. On the other hand, a Bayesian extension of ODIL could potentially provide quantified uncertainties for inverse problems in higher dimensions, as ODIL was shown to be orders of magnitude faster and more robust than PINNs in two and three dimensional benchmark problems [30].
In this study, we present B-ODIL, a Bayesian extension of ODIL. B-ODIL enables uncertainty quantification and robustness to model error in high-dimensional PDE-based inverse problems, while retaining the computational structure and scalability of ODIL. In this framework, the prior incorporates knowledge from the PDE, while the likelihood couples the observed data to the unknown field. This method provides a solution to the inverse problem with quantified uncertainties. We account for the computational burden by estimating the posterior distribution with different sampling techniques and two approximation strategies: (i) a full Laplace approximation to the joint posterior for moderate-dimensional problems, and (ii) a parameter-focused mode approximation designed for scalable inference in large-scale cases. We have designed a series of benchmarks with increasing levels of complexity to test B-ODIL. First, we consider the ODE describing the dynamics of a harmonic oscillator, to compare the validity of the Laplace and mode approximations compared with Hamiltonian Monte-Carlo (HMC) sampling of the same posterior in a computationally tractable setting. We then apply B-ODIL to the one dimensional PDE of the diffusion equation with unknown initial conditions, introducing the challenge of ill-posed inverse problems, typical in PDE-based inference, where uncertainty quantification is crucial. The third benchmark tests the method’s ability to reconstruct the states of a non-linear two dimensional PDE from synthetic data, and we demonstrate that the ground truth falls within the uncertainty bounds provided by B-ODIL. Finally, we apply the method to a three dimensional model of tumor growth coupled with real patients data, and provide estimates of tumor cell fields with quantified uncertainties given medical images.
@@ 2.1 ODIL
where $\lambda$ is a positive constant that describes the importance of fitting the data with respect to satisfying the discrete PDE. In this work, the minimization of eq. 4 is performed with gradient-based methods implemented in Pytorch [44] and detailed in the results section.
@@ 2.2 B-ODIL: Bayesian inference for inverse problems with ODIL
We now consider the Bayesian formulation of PDE-based inverse problems, where the goal is to infer parameters $\theta$ and the solution field $\mathbf{u}$ from noisy measurements $\mathcal{D}=\{y_{k}\}_{k=1}^{N_{D}}$, given a model in the form of a PDE. In this framework, the posterior distribution is obtained by combining a likelihood model for the data with prior knowledge enforcing compatibility between the solution field and the PDE. As we have seen in the previous section, ODIL provides a natural way to encode this compatibility through a PDE loss. We thus incorporate this PDE loss into the prior information, allowing us to extend ODIL into the Bayesian setting. In classical Bayesian inverse problems with well-posed PDEs and accurate forward models, one often eliminates the state variable via a map $\mathbf{u}=f(\mathbf{\theta})$, yielding a reduced posterior over $\mathbf{\theta}$. In contrast, B-ODIL targets settings in which the governing equations are enforced approximately through a residual-based term and may be affected by discretization error or model misspecification. In this regime, enforcing a strict parameter-to-state map can lead to biased and overconfident posteriors, motivating the use of a joint posterior over $(\mathbf{u},\mathbf{\theta})$ that can accommodate structural model error. We assume that the measurements $\mathcal{D}$ are noisy and that we have a model describing this noise. We would like to estimate the parameters $\mathbf{\theta}$ and the solution $\mathbf{u}$, with quantified uncertainties. According to the Bayes’ theorem, the posterior distribution of these quantities given the data reads
@@
where $\sigma$ is the standard deviation associated with the observable and $h_{k}$ , $k=1,\dots,N_{D}$, are measurement operators on the solution $\mathbf{u}$. For example, $h_{k}$ can be a linear interpolation of the field at the measurement location from the discrete field $\mathbf{u}$. Throughout this work, measurements are assumed conditionally independent given $(\mathbf{u},\mathbf{\theta})$, corresponding to independent observation noise; correlations between measurements could be incorporated through a non-diagonal likelihood covariance but are not considered here for simplicity. Other likelihood models can be used within the same framework. Examples are provided in the results section.
As prior knowledge, we assume that the possible solutions $\mathbf{u}$ are compatible with the PDE associated with the model. To account for this knowledge, the prior over the solution $\mathbf{u}$ with parameters $\mathbf{\theta}$ is such that the loss $L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})$ is small, and thus we choose priors of the form
$P(\mathbf{u},\mathbf{\theta})=\frac{1}{Z}\exp\left(-\beta L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})\right)P(\mathbf{\theta}),$ (7)
where $\beta$ is a positive scalar that controls how peaked the distribution is, and $Z$ is a normalization constant that does not depend on $\mathbf{u}$ or $\mathbf{\theta}$. Finally, $P(\mathbf{\theta})$ contains additional prior knowledge on $\mathbf{\theta}$. In this work, $P(\mathbf{\theta})$ is taken as improper uniform prior in all cases. The prior in eq. 7 is defined at the discrete level through the loss $L_{\mathrm{PDE}}$. As a result, the posterior is a discretization-dependent object. In this work we adopt a computational perspective and evaluate posterior stability under mesh refinement (see section 3.1), which provides an empirical check that discretization effects are negligible once posterior summaries have converged. We refer to Stuart [49], Cotter et al. [14] for the function-space formulation of Bayesian inverse problems and discretization-robust algorithms. The posterior distribution of the solution and parameters of the model thus becomes
$P(\mathbf{u},\mathbf{\theta}|\mathcal{D})\propto P(\mathcal{D}|\mathbf{u},\mathbf{\theta})\frac{1}{Z}\exp\left(-\beta L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})\right)P(\mathbf{\theta}).$ (8)
The dimensionality of this problem is large due to the unknown variables $\mathbf{u}$, challenging sampling techniques. To obtain tractable uncertainty estimates in high dimensions, we use the Laplace approximation to the joint posterior. The log-posterior reads
$\log P(\mathbf{u},\mathbf{\theta}|\mathcal{D})=\log P(\mathcal{D}|\mathbf{u},\mathbf{\theta})-\beta L_{\mathrm{PDE}}(\mathbf{u},\mathbf{\theta})-\log{Z}+\log P(\mathbf{\theta})+C,$ (9)
where $C$ is a scalar that does not depend on $\mathbf{\theta}$ or $\mathbf{u}$. We expand this quantity in Taylor series up to second order around the maximum a posteriori (MAP), where the linear term vanishes, leading to the Laplace approximation [53]
@@
where $\mathbf{v}=(\mathbf{u},\mathbf{\theta})$, $\mathbf{v}^{\star}$ is the solution at the MAP, and $H$ is the Hessian of the log-posterior evaluated at $\mathbf{v}^{\star}$. Thus, the posterior is approximated with a multivariate Gaussian with mean $\mathbf{v}^{\star}$ and covariance $\Sigma=-H^{-1}$. As with any local Gaussian approximation, the Laplace method is expected to be inaccurate for strongly non-Gaussian or multimodal posteriors. In such regimes, full sampling or alternative approximations would be required. We note that when the likelihood takes the form of eq. 6, maximizing eq. 9, i.e. finding the MAP, corresponds to the original ODIL method, eq. 4 with $\lambda=N_{D}/\beta$, assuming that $P(\mathbf{u})$ and $P(\mathbf{\theta})$ are taken as improper uniform priors. Thus, computing uncertainties over predictions that were computed with ODIL consists only in computing the Hessian and inverting it. In general, forming and inverting the Hessian with respect to $\mathbf{v}=(\mathbf{u},\mathbf{\theta})$ is computationally prohibitive when $\mathbf{u}$ is high-dimensional; consequently, the full Laplace approximation is feasible only for relatively low-dimensional benchmarks, while large-scale problems require parameter-focused or reduced approximations discussed below.
@@ 2.3 Inference of model parameters for inverse problems
The evaluation of this high-dimensional integral is intractable. Instead, we assume that the joint distribution is peaked around the MAP $\mathbf{u}^{\star}(\mathbf{\theta})=\operatorname*{arg\,max}\limits_{\mathbf{u}}P(\mathbf{u}|\mathbf{\theta},\mathcal{D})$ and we approximate the posterior distribution of the model parameters using the mode approximation [53]
@@
The formulation adopted here is not intended as a replacement for classical Bayesian inference based on repeated forward model evaluations under a strict parameter-to-state map $\mathbf{u}=f(\mathbf{\theta})$. Such reduced formulations are appropriate when the forward model is accurate and can be evaluated efficiently for each value of $\mathbf{\theta}$. In contrast, B-ODIL targets settings in which the governing equations may be misspecified. We note that several studies have approached this problem by directly altering the model [41] or by statistically correcting its output [55]. Here, we instead incorporate the PDE residual as a soft probabilistic constraint in the prior. Embedding optimization over $\mathbf{u}$ within inference over $\mathbf{\theta}$ allows the PDE constraints and the data to be reconciled at the field level, while the parameter $\beta$ controls tolerance to model error. As shown in section 3.2, this is essential for obtaining calibrated posteriors under misspecification, where standard Bayesian inference can otherwise become overconfident and biased.
2.4 Selection of the parameter $\beta$
The parameter $\beta$ controls the relative importance of the PDE-residual prior and the data likelihood. When the governing equations are known to be accurate, large values of $\beta$ enforce strong adherence to the model. The limit $\beta\rightarrow\infty$ corresponds to classical methods for PDE-based Bayesian inverse problems. In the presence of model error or misspecification, however, overly large $\beta$ can lead to biased and overconfident posterior estimates.
In practice, $\beta$ can be selected in different ways depending on the availability of data and computational resources. A principled approach consists in selecting $\beta$ by maximizing the predictive log-likelihood on held-out data. Specifically, the available observations are split into training and validation sets. For each candidate value of $\beta$, B-ODIL is applied to the training data to obtain a posterior distribution, which is then evaluated on the validation data. The value $\beta^{\star}$ maximizing the validation log-likelihood is retained.
When validation is impractical (e.g., limited data or high computational cost), we select $\beta$ by tuning, guided by the same trade-off as in classical ODIL and by posterior diagnostics (e.g., predictive uncertainty and residual structure). Alternatively, $\beta$ can be treated as a hyperparameter and selected by maximizing the model evidence, at the cost of introducing a prior on $\beta$ and additional approximation or optimization steps. In the present work we primarily use cross-validation when feasible, and otherwise adopt pragmatic tuning that is consistent with ODIL practice and supported by posterior predictive checks.
@@ 3.1 Harmonic oscillator
We first consider a simple example to demonstrate the validity of the Laplace approximation and that of eq. 12. Consider the system of ODEs of a harmonic oscillator with position $x$, velocity $v$, mass $m$ and spring coefficient $k$,
@@
for time $t\in(0,T)$ and with unknown initial conditions. Given measurements of position at known times $\mathcal{D}=\left\{t_{j},x_{j}\right\}_{j=1}^{N_{D}}$ with known standard deviation $\sigma$, we want to predict the position and velocity of the system throughout the whole time interval. We discretize the position and velocity into $N=64$ equidistant time intervals with $\Delta t=T/N$ and $T=20$. The discrete losses are computed using the midpoint rule,
@@
Finally, the observation functions correspond to linear interpolations of the positions at data points $t_{j}$, $j=1,2,\dots,N_{D}$. For now, we assume that the parameter $\omega=\sqrt{k/m}=1$ is known. We use synthetic data consisting of $N_{D}=20$ time points uniformly distributed in $t\in(0,T/2)$, and corresponding positions with Gaussian distribution centered around the exact solution of the model with standard deviation of $\sigma=0.1$. The initial conditions used to generate the data were set to $x(0)=0.5$ and $v(0)=0.2$. We choose a value of $\beta=${10}^{4}\text{\,}$\$ in the prior distribution (see eq. 7). The MAP is computed by maximizing the posterior with respect to $(\mathbf{x},\mathbf{v})$ with the Adam optimizer [31] with learning rate ${10}^{-3}\text{\,}$ for 10’000 epochs. Computing the Laplace approximation for this case takes about $2\text{\,}\mathrm{s}$ on a consumer-grade laptop. The Laplace approximation of this problem is shown on fig. 1. For comparison, we solve the same problem with HMC on the same posterior, and show the results on the same figure. The HMC results were obtained from a custom implementation of the original HMC algorithm [43] with 10’000 samples, chain lengths $l=10$, mass matrix $M=1$, and step size $\delta t=0.008$, tuned for an acceptance rate of $\alpha\approx 0.65$. This implementation with these parameters took about $5\text{\,}\mathrm{min}$ on a consumer-grade laptop.
Figure 1: Left: Prediction of the position (top row) and velocity (bottom row) of the oscillator given the data (crosses) using the B-ODIL framework with Laplace approximation (left column) and HMC (right column). The shaded area denotes the 5 to 95% quantiles of the posterior, and the solid line denotes the posterior mean. The dashed line represents the underlying process that was used to generate the data. Right: Covariance matrices of the full solution $(x_{1},\dots,x_{N},v_{1},\dots,v_{N})$ obtained from the Laplace approximation (left column) and estimated from HMC samples (right column).
@@
$P(\mathbf{\theta}|\mathcal{D})=\frac{(2\pi)^{d/2}}{\sqrt{\det H(\mathbf{\theta})}}\exp l(\mathbf{u}^{\star}(\mathbf{\theta}),\mathbf{\theta}),$
@@
where $d$ is the dimension of $\mathbf{u}$. To evaluate this expression for each value of $\theta$, we use the optimal value $\mathbf{u}^{\star}(\mathbf{\theta})$ estimated with ODIL, and use automatic differentiation in Pytorch to compute $H(\mathbf{\theta})$. We use this approach throughout this work. In practice, $\mathbf{u}^{\star}(\mathbf{\theta})$ is obtained by numerical optimization and is therefore only approximate; the resulting expression should be interpreted as a Laplace-type approximation based on a locally quadratic expansion around a numerically converged solution.
Figure 3: Left: Prediction of the position $x$ at time $t=20$ with the Laplace approximation against the grid size $n_{t}$. Symbols represent the mean prediction and shaded areas represent the 50% and 90% confidence intervals. Right: Variance of the prediction of $x$ at time $t=20$ against the parameter $\beta$, with $N_{D}=10,100$, and $1000$, from top to bottom. The solid black line shows a linear decrease with $\beta$.
Finally, we study the effect of the grid size $N$ and parameter $\beta$. First we fix the PDE weight $\beta=${10}^{4}\text{\,}$\$ and the number of measurements $N_{D}=20$, and predict the position $x$ at the last grid point ($t=20$) given the data and the discretized PDE. Figure 3 shows that the posterior of $x(t=20)$ converges as we increase the grid size. In practice, and throughout this work, we select the smallest grid resolution for which posterior summaries have converged, in order to minimize computational cost while avoiding discretization-induced artifacts.
Second, we vary the parameter $\beta$ for three values of $N_{D}$ and with a fixed resolution $N=256$, for which the posterior has converged. The variance of the posterior prediction of $x$ at time $t=20$ is shown against $\beta$ in fig. 3. For each value of $N_{D}$, we observe two regimes. For low values of $\beta$, corresponding to soft priors on the model, the variance decreases linearly with $\beta$. When $\beta$ is large, the variance reaches a plateau that is limited by the uncertainties coming from the data term. As expected, a larger amount of data $N_{D}$ decreases the value of this plateau. In the limit $\beta\rightarrow\infty$, the discretized PDE is enforced exactly, and thus the only uncertainties come from the data. Empirically, larger $\beta$ values increase the conditioning of the optimization problem and the number of iterations required for convergence, without improving predictions once the posterior variance has reached the data-limited plateau. We thus use values of $\beta$ that are smaller than the start of the plateau. From a Bayesian perspective, $\beta$ controls the tolerance to model error by tempering the influence of the PDE-based prior relative to the data likelihood. In many practical cases, we do not know models that are perfectly consistent with observations, in which case a lower value of $\beta$ is favored. We give more details about this choice in the next section.
3.2 Choice of the parameter $\beta$ under model misspecification
We now illustrate the practical impact of the $\beta$-selection strategy introduced in section 2.4 in a setting with deliberate model misspecification. Specifically, we consider a variation of the harmonic oscillator example from section 3.1, in which the data are generated by an unknown nonlinear oscillator while inference is performed using the linear model. The nonlinear oscillator follows the dynamics
$\frac{dx}{dt}$ $=v,$ (16)
$\frac{dv}{dt}$ $=-\frac{1}{m}(k_{1}x+k_{2}x^{3}),$ (17)
where $k_{1}$ and $k_{2}$ are positive constants. We generate $N_{D}=200$ data points using $x_{0}=2$, $v_{0}=0$, $m=15$, $k_{1}=1$, $k_{2}=2$, and noise $\sigma=0.4$. We then apply the B-ODIL framework with the linear model described by eq. 15, using the Adam optimizer with a learning rate of $5\text{\times}{10}^{-4}\text{\,}$ over 50000 epochs. This procedure mimics many practical cases where the statistical model is an approximation of the true underlying data generation process. We split the data into training and validation sets (80% and 20% of the data, respectively), run B-ODIL on the training set with different values of $\beta$, and select $\beta^{\star}$ by maximizing the validation log-likelihood.
Figure 4: Predictions of and oscillator dynamics with a linear spring model with data coming from a nonlinear spring. Left: log-likelihood of the validation set against the parameter $\beta$. We select the value $\beta^{\star}$ that maximizes this log-likelihood. Right: Posterior prediction of the position and velocity field with the B-ODIL framework with $\beta=\beta^{\star}$ (left column), and with conventional Bayesian inference (right column). The solid black line shows the exact underlying process used to produce the data.
Figure 4 shows the log-likelihood of the validation set against $\beta$. High values of $\beta$ result in a plateauing log-likelihood, as the method tends to conventional Bayesian inference as $\beta\rightarrow\infty$. On the other hand, low values of $\beta$ lead to a decrease of the log-likelihood as the model is less and less trusted to make sharp predictions. This log-likelihood has a maximum around $\beta^{\star}\approx 5.7$, which balances fitting the data and trusting the model. We use this value to make predictions and show the posterior distribution of the position and the velocity in fig. 4. The 90% confidence intervals of the posterior contain the true underlying process. In contrast, conventional Bayesian inference based on the misspecified linear model produces overly concentrated posteriors that fail to account for structural model error, leading to confidence intervals that do not capture the true dynamics away from the data (fig. 4). This result was obtained with the MCMC sampling method with 100’000 samples and a proposal step size tuned to have a 35% acceptance rate.